16 factory workers made 40 pieces in a certain amount of time.
a.Do these quantities (numer of workers and number of pieces produced) vary directly or inversely?
b.Find the constant of variation.
c.How many workers are needed to make (1) 20, (2)25 and (3)100 pieces in the same amount of time, assuming each worker works at the same rate?

Answer by Boreal(15235)

directly because the more workers the more number of pieces produced
16=40*k
k=0.4
W=0.4*20=8 workers
W=0.4*25=10 workers
W=0.4*100=25 workers